OpenStudy (anonymous):

Determine the value of C that satisfies this equation. Let u=-i+2j+3k. IIcuII=7. where C is a scalar

OpenStudy (anonymous):

||u|| = sqrt( (-1)^2 + (2)^2 + (3)^2 ) = sqrt( 1+ 4 + 9) = sqrt(14) so c is the number which you multiply it by sqrt(14) and get 7

OpenStudy (anonymous):

the answer at the back of the book says plus or minus 7/3. The answer you gave is similar to what I got. I got 7/sqrt14

OpenStudy (turingtest):

if the answer was 7/3 that would lead to\[\sqrt{(\frac73)^2+4(\frac73)^2+9(\frac73)^2}=\frac73\sqrt{14}\neq7\]so your book is screwy, the answer is yours 7/sqrt14, which simplifies to sqrt(7/2)

OpenStudy (anonymous):

thanks

OpenStudy (turingtest):

welcome, don't trust text books :)